You were told by your teacher to makeva sketch of a rectangular garden such that is perimeter is at 19cm of the difference of length and width least 5m

Found 2 solutions by ikleyn, macston:
Answer by ikleyn(52781)

.
You were told by your teacher to makeva sketch of a rectangular garden such that is perimeter is at 19cm of the difference of length and width least 5m
-------------------------------------------------------------------

It was you "were told", not me.

Check your tone when speak to your tutor.


Answer by macston(5194)

.
If you mean the difference between length and width
is at least 5 cm (not 5 meters as you posted), then:
.
L=length; W=width
.
P=2(L+W)
19=2(L+W)
9.5=L+W
.
L-W=5
L=5+W
.
9.5=L+W
9.5=(5+W)+W
4.5=2W
2.25=W
.
L=W+5=2.25+5=7.75
.
ANSWER: One solution is a 7.75cm x 2.25 cm rectangle.